Please help ,my son wants to quit school over this and I have no idea haw to help him.I don't have money to pay for private help .
Ne 1
Solve the following system of equations for X and Y .Elimination , keep as fractions .
1 + 3 = 3
--- --- ----
2X + 2y = 8

9 - 6 = 15
-- --- ----
x - y = 12
Nr 2 solve the following system of equations for x,y,z
2x - 3y + z = 10
3x + 6 z = 24
- 5 y - 20z = -55 substitution
nr 3
A triangle is formed by the intersection of the lines 2x + 3y = 14
4X - 5y= - 16 and the x- axis.Find the area of the triangle

Answer by ReadingBoosters(3246)

No. 1
Starting with
1/2x + 3/2y = 3/8
9/x - 6/y = 15/12
Multiply the first equation by 6 and the second equation by 3/2
6((1/2x) + (3/2y) = (3/8))
3/2((9/x) - (6/y) = (15/12))
to yield
6/2x + 18/2y = 18/8
27/2x - 18/2y = 45/24
Add the equations together
6/2x + 27/2x + 18/2y + -18/2y = 18/8 + 45/24
Combine like terms
33/2x = 18/8 + 45/24
Multiply both sides by 2x
33 = 2x(18/8 + 45/24)
33 = 36x/8 + 90x/24
Give the left side common denominators
33 = 36x/8(3/3) + 90x/24
33 = 108x/24 + 90x/24
33 = 108x+90x/24
33 = 198x/24
Multiply both sides by 24
792 = 198x
x = 4
...
Starting with
1/2x + 3/2y = 3/8
9/x - 6/y = 15/12
Multiply the first equation by -9 and the second equation by 1/2
-9(1/2x + 3/2y = 3/8)
1/2(9/x - 6/y = 15/12)
to yield
-9/2x - 27/2y = -27/8
9/2x - 6/2y = 15/24
Add the equations together
-9/2x + 9/2x + -27/2y + -6/2y = -27/8 + 15/24
Combine like terms
-27/2y - 6/2y = -27/8 + 15/24
-33/2y = -27/8 + 15/24
Give common denominators on the right side
-33/2y = -27/8(3/3) + 15/24
-33/2y = -81/24 + 15/24
-33/2y = -66/24
Multiply both sides by 2y
-33 = 2y(-66/24)
-33 = -132y/24
Multiply both sides by 24
-792 = -132y
Divide both sides by -132
y = 6
...
x = 4, y = 6
...
No. 2
Starting with
2x - 3y + z = 10
3x + 6z = 24
-5y - 20z = -55
Rearrange the second equation
3x+6z=24 to x = (24-6z)/3
Rearrange the third equation
-5y-20z=-55 to y = (55-20z)/5
Plug x and y into the first equation to solve for z
2((24-6z)/3) - 3((55-20z)/5) + z = 10
2(24/3 - 6z/3) - 3(55/5 - 20z/5) + z = 10
2(8 - 2z) - 3(11 - 4z) + z = 10
16 - 4z - 33 + 12z + z = 10
Combine like terms
9z - 17 = 10
9z = 27
z = 3
...
Plug z into -5y - 20z = -55
-5y - 20(3) = -55
-5y - 60 = -55
-5y = 5
y = -1
...
Plug z into 3x + 6z = 24
3x + 6(3) = 24
3x + 18 = 24
3x = 6
x = 2
...
Confirm using 2x - 3y + z = 10
2(2) - 3(-1) + 3 = 4 + 3 + 3 = 10
...
No. 3
Repost No. 3 separately
............................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.